OpenStudy (anonymous):

3x^2-x-10=0 What is the value for c?

OpenStudy (akashdeepdeb):

3x^2-x-10=0 3x^2 - (6-5)x - 10 =0 3x^2 - 6x + 5x - 10 = 0 3x(x-2) + 5(x-2) (3x+5)(x-2) = 0 So x can be -5/3 or +2 Understood? :)

OpenStudy (wolfe8):

First of all I don't understand by the value of c. @AkashdeepDeb I think what you calculated is the root of x. Please explain to me what c is. I might have forgotten about this.

OpenStudy (akashdeepdeb):

I think it Is x ad NOT c :)